OpenStudy (anonymous):

A force of 6.00 newtons on a car's bumper makes an angle of 28.5 degrees with the ground. What is the vertical component of this force vector?

OpenStudy (ghazi):

 the vertical component of force is given by \[F _{V}=F \sin \theta= 6* \sin 28.5 N\]